your quiz grades are 78,68,71,85, and 90. what score on the sixth quiz will make your average quiz at least 80? what is the equation and solve the problem?

Answer by checkley79(3341)

(78+68+71+85+90+X)/6>=80
(392+X)/6>=80 CROSS MULTIPLY
392+X=6*80
392+X=480
X=480-392
X=88 ANS, FOR THE MINIMUM SCORE TO MAINTAIN AN 80 AVERAGE OR GREATER.
PROOF:
(392+88)/6=80
480/6=80
480=6*80
480=480